# Approximating a gene regulatory network from non-sequential data

### v2

@@ 3 Problem statement and overview
where $W_{t}$ is $n$-dimensional standard Brownian motion, $A$ is the diffusivity matrix, $AA^{T}$ is the covariance matrix of increments, and $B$ the drift matrix.
Suppose that we observe $x_{t}$ at various time points in $[0,T]$ for large $T$, but we either do not have the ordering information or do not keep the samples ordered. Thus the covariance matrix of the increments of $x_{t}$ is not assumed to be known in this paper, since it requires ordering information for its computation. Then, we want to know whether we can recover the underlying (unknown) $n\times n$ drift matrix $B$, that characterizes our dynamical system. [^2] We have assumed the system is centered about 00 but that is not an issue since the long term average can be computed without knowing the ordering and subtracted from the observations.
Remark 3.2.
If we are given an ordering of $x_{t}$s (correct or with errors) as $(x_{t_{0}},x_{t_{1}},...,x_{t_{n}})$ then we may compute covariance matrix of increments, with respect to that ordering, as the average $\frac{1}{n}\sum_{k=1}^{n}(x_{t_{k}}-x_{t_{k-1}})^{T}(x_{t_{k}}-x_{t_{k-1}})$. If $t_{0}\leq t_{1}..\leq t_{n}$ then $AA^{T}$ is the empirical covariance matrix of increments and is closely approximated by the average, by Sanov type concentration theorems (See Dembo and Zeitouni (2010)). Moreover, being an average, this matrix is fairly robust to noise and small errors; a fact that we will use critically in this paper.
In general, reconstructing $B$ without ordering information of the $x_{t}$ may not be possible. However, the main technical result in this paper shows that if: (1) there exists an ”approximate clock” (see Definition 3.3) and (2) $B$ is row sparse, i.e., most of the entries in each row of $B$ are zero or close to zero; then we can approximately reconstruct $B$ with few samples in polynomial time, by solving a convex program.
Definition 3.3.
@@
To keep the presentation simple, we will assume that if there exists a single coordinate of $x(t)$, [^4] Here we assume a single coordinate exists as opposed to multiple coordinates or allowing a linear combination of coordinates denoted by $\tau(t)$, and it is modeled by the diffusion:
@@
where $W^{\prime}_{t}$ is standard Brownian motion and $\delta$ is positive and much greater than $\varepsilon^{2}$; then $x(t)$ is said to admit an approximate clock. [^5] Notation: We use $x_{t}$ and $x(t)$ interchangeably.
@@
We will use the approximate clock to recover an approximation of the matrix $A$, which we denote by $\tilde{A}$. We won’t actually need $\tilde{A}$ but rather the covariance of the increments: $\tilde{A}\tilde{A}^{T}$; and in Theorem 4.1 in Section 4, we show that ordering the $x_{i}$ values using $\tau(t)$ and computing the covariance from the increments closely approximates the covariance of the increments computed using the ordering based on the actual time.
Algorithm 1 runs in polynomial time as the main steps are computing a covariance matrix and solving the convex optimization problem in Step 7 of Algorithm 1, which is a second order cone program. This convex optimization is a slight variation of basis pursuit in the compressed sensing literature (see for example Foucart and Rauhut (2013)) and therefore unique reconstruction is guaranteed under sparsity assumptions. In particular the following uniqueness theorem holds true for the matrix $\tilde{B}$ recovered by Algorithm 1.
Theorem 3.4.
Dirksen et al. (2018) Let $s$ denote the row sparsity of $B$, i.e., each row has at most $s$ non-zero coordinates, then for $m=\Omega(\log n)$ large and the sample times $\{t_{1},...,t_{m}\}$ for samples $\{x_{t_{1}},...,x_{t_{m}}\}$ are uniformly distributed, then with probability $1-o(1)$ (for large $n$): $\|\tilde{B}-B\|_{1}=0$, i.e., $B$ can be recovered uniquely by solving a convex program. If however, the rows are not exactly $s$ sparse and the $\ell_{1}$ norm of the $n-s$ remaining entries in any row is at most $\epsilon$, then we have: $\|\tilde{B}-B\|_{1}\leq\epsilon n$.
While the proof follows by easy modifications of known results in existing compressed sensing literature (see Foucart and Rauhut (2013) or Dirksen et al. (2018)), we note the one novelty here over existing compressed sensing literature: the formulation of the constraint in Step 7 of Algorithm 1 uses the characterization of the long term variance of the OU process (see Oksendal (2013)).
Algorithm 1 recovers the drift matrix $B$ from the covariance matrix $\Sigma$ and the (approximate) covariance matrix of the increments $\tilde{A}\tilde{A}^{T}$ up to an additive factor (that is just the ratio of diffusivity over drift of the approximate clock).
@@
5:  $\triangleright$ Note:(1) computing $\Sigma$ does not require ordering the $x_{i}$s
6:  $\triangleright$        (2) $B(i,\cdot)$ denotes the $i^{th}$ row of $B$
7:  Solve $\tilde{B}:=\arg\min_{B}\sum_{i=0}^{n}\|B(i,\cdot)\|_{2}$     s.t. $\Sigma B+B^{T}\Sigma=-\frac{\tilde{A}\tilde{A}^{T}}{2}$
8:  return $\tilde{B}$
Remark 3.5.
@@
5:  $\triangleright$ The smaller number of samples prevents direct computation of $\tilde{B^{\prime}}$, so we compute a first order approximation.
@@
7:  $\triangleright$ Choose a small noise $\eta$ so that the following convex program program is feasible
8:  $\triangleright$ Note: $P_{ij}$ denotes the $i^{th}$ row, $j^{th}$ column element of matrix $P$
@@
3.1 Algorithm for isolating genomic pathways
@@
  1. 1.
    If the gene corresponding to the terminal node of $p$ positively correlates with the pathological phenotype then it should be up-regulated by the gene preceding it in the path $p$; moreover, the same property or property (2) should now hold for the subpath terminating in the penultimate gene, and so on.
  2. 2.
    Otherwise, if the gene corresponding to the terminal node of $p$ negatively correlates with the pathological phenotype then it should be down-regulated by the gene preceding it in the path $p$; moreover, the same property or property (1) should now hold for the subpath terminating in the penultimate gene, and so on.
@@
3:  $\triangleright$ Algorithm starts:
4:  Compute and sort the list of genes in descending order in terms of the absolute value of their correlation coefficient with the pathological phenotype, denote this list as $\vec{g}$
5:  Compute $C:=\mathrm{Diag}(\vec{g})(\tilde{B^{\prime}}-\tilde{B})$
@@
The results of running Algorithm 3 on the breast cancer dataset (Cancer Genome Atlas Network: Daniel C Koboldt et al, 2012) to isolate critical differentiating pathways are described in Figure 1. In this figure, we show predicted pathways where gene A $\to$ gene B means that A (indirectly) up-regulates B, while the flat arrowhead indicates (indirect) down-regulation. A green node indicates that the corresponding gene expression positively correlates with the pathological phenotype (metastasis), while a red node indicates positive correlation with the reference phenotype (cancer without metastasis). We predict potential pathways, restricted to nuclear receptors, for metastasis. Based on Figure 1(a), one prominent pathway could be: NR1I2 up-regulates RORB which up-regulates RORC. A simple correlation coefficient analysis of RNA-seq data in Cancer Genome Atlas Network: Daniel C Koboldt et al (2012) shows that the expression of RORC is positively correlated with metastasis. However, the actual role of RORC in metastasis is likely to be far more complicated (see Oh et al. (2016)).
4 Theoretical bounds on error from an approximate clock
In this section, we derive an upper-bound on the error of the recovered diffusivity matrix when we use an approximate clock. We will assume that our dynamical system can be modeled by an OU process and it has a coordinate, or a linear combination of coordinates, that admit a positive drift which is larger than the diffusivity. This coordinate will act as an approximate clock. Theorem 4.1 shows that if we use this approximate clock then the recovered diffusivity is not too different from the diffusivity that would be obtained if we knew the ordering in the dynamical system.
Theorem 4.1.
Assuming that the clock coordinate $\tau(t)$ is independent of the remaining coordinates of $x_{t}$, the diffusivity matrix $\tilde{A}\tilde{A}^{T}$ computed using the approximate clock $\tau(t)$ is asymptotically equal to the actual diffusivity, i.e.,
$\tilde{A}\tilde{A}^{T}=\lim_{h\to 0}\mathbb{E}\left[\frac{1}{\delta T}\int_{0}^{T}\langle(x(\tau(t+h))-x(\tau(t))),\right.\\ \left.(x(\tau(t+h))-x(\tau(t)))\rangle d\tau(t)\right]\simeq AA^{T},$ (3)
for $\varepsilon\ll\delta$ [^6] Note: (1) Since $\mathbb{E}[\tau(t)]=\delta t$ the normalization of $\tilde{A}\tilde{A}^{T}$ has a $\delta T$ term as opposed to a $T$ term, (2) one can skirt around the issue of negative values of $\tau$ by assuming that $\tau(0)$ is large, that will ensure $\tau$ remains positive with high probability.
The proof is deferred to the supplement (Section A.1). If the correlations between the randomness in $\tau(t)$ and other coordinates are significant then we can evaluate Equation 12 as follows.
Lemma 4.2.
Let $\rho\in\mathbb{R}^{1\times n}$ be the covariance matrix between $W^{\prime}_{t}$ and $W_{t}$, which are the Brownian motions in the definitions of $\tau(t)$ and $x_{t}$ respectively; then the quadratic variation below evaluates as:
$\mathbb{E}\left[\langle AdW_{\tau},Bx(\tau)d\tau\rangle\right]=\varepsilon A\rho x(t)^{T}B^{T}dt\ .$ (4)
The proof follows immediately from the definition of quadratic variation. Therefore, following a similar proof as Theorem 4.1, the error to $\tilde{A}\tilde{A}^{T}$ from Theorem 4.1 is changed by an additive factor that is $O(\varepsilon\max_{i.j}\{\rho_{ij}\})$. So when all coordinates of $\varepsilon\rho$ are small, the characterization of $\tilde{A}\tilde{A}^{T}$ in Theorem 4.1 continues to hold.
5 Empirical bounds on error from an approximate clock
Theorem 4.1 shows that the covariance matrix $AA^{T}$ in Algorithm 1 is determined by an approximate clock as long as $\frac{\epsilon}{\delta}$ is small. In this section we verify the same fact empirically.
5.1 Empirical verifation using synthetic data
In order to generate the plot in Figure 2, we used an OU process with a $10\times 10$ identity matrix as a stand-in for $A$; we set $B:=-Id+G$, where $G$ is a random Gaussian matrix, and we used time-steps of size $0.0001$ to generate the sample path for $1000$ time-steps. We introduce $k$ errors/inversions (the $x$ coordinate in Figure 2) in the sample path by picking $k$ pairs of points, uniformly at random, in the sample path and swapping the pairs. Finally, we compute the corresponding empirical covariance matrix: $\tilde{A}(k)\tilde{A}(k)^{T}$, from the perturbed path with $k$ inversions, the corresponding Frobenius norm of $\tilde{A}(k)\tilde{A}(k)^{T}-AA^{T}$ and plot the latter as a function of $k$.
5.2 The approximate clock using single cell RNA-seq datasets
(a) Dynamical system obtained from latent time computation
(b) Dynamical system obtained from Cyclin clock
(d) Correlations between cyclin expressions and scVelo latent time
It can be argued that neither Theorem 4.1 nor Figure 2 uses cyclin expressions anywhere. Perhaps using cyclin expressions to order cells leads to a significantly large number of errors in cell sample orderings of RNA-seq data. In this subsection, we allay that fear as well.
Yet another way to verify that the approximate clock using cyclin expressions does not lead to a large error in-between $AA^{T}$ and $\tilde{A}\tilde{A}^{T}$ is to use single cell RNA-seq data. Such datasets admit an ordering of cell samples using latent time based on underlying RNA velocity models. In this subsection, we verify that $AA^{T}$ (determined using latent time) and $\tilde{A}\tilde{A}^{T}$ (determined using a simple ordering using cyclin expressions) are close, based on single cell RNA-seq dataset from murine pancreatic tissue (Bastidas-Ponce et al., 2019).
So we compute the covariance matrix of increments, among significantly expressed nuclear receptors in the murine pancreatic scRNA-seq dataset from (Bastidas-Ponce et al., 2019), in two ways:
@@ 5.2 The approximate clock using single cell RNA-seq datasets
    First, we compute the matrix $\tilde{A}\tilde{A}^{T}$ using an approximate clock that involves ordering the cell samples by the difference in between two cyclin expressions, say Cyclins I and D, i.e., $\tau(c):=\mathrm{Cyclin\ I}-\mathrm{Cyclin\ D}$ for a given cell $c$. Cyclin I and D represent the optimal choice based on the correlations in Figure 3(d).
@@
    Next, we compute the matrix $AA^{T}$ by using the latent time function in scVelo as our ordering (this is our ground truth).
We use the total RNA expression levels data for nuclear receptors in the five groups: alpha, beta, epsilon, pre-endocrine and ductal cells in the murine pancreatic single cell RNA-seq dataset (Bastidas-Ponce et al., 2019) for this purpose. Thus we obtain five instances of $AA^{T}$ and five instances of $\tilde{A}\tilde{A}^{T}$.
We performed a similar analysis on the murine dentate gyrus dataset of Hochgerner et al. (2018). The covariance matrices computed in two ways (using scVelo latent time and cyclin expression based ordering) are shown in Figure 5 in the supplement.
6 The reconstructed dynamical system from breast cancer data
Figure 4: Perturbations of GRN towards metastasis. The matrix represents the first order approximation of changes in interaction between genes in presence of metastasis. Values far from 00 (yellow) represent a prediction that there is significant up or down regulation in an interaction between a given pair of genes.

### v3

@@ 2 Technical contributions
where $W_{t}$ is $n$-dimensional standard Brownian motion, $A$ is the diffusivity matrix, $AA^{T}$ is the covariance matrix of increments, and $B$ the drift matrix. (see for example oks)
Suppose that we observe $x_{t}$ at various time points in $[0,T]$ for large $T$, but we either do not have the ordering information or do not keep the samples ordered. Thus the covariance matrix of the increments of $x_{t}$ is not assumed to be known in this paper, since it requires ordering information for its computation. Then, we want to know whether we can recover the underlying (unknown) $n\times n$ drift matrix $B$, that characterizes our dynamical system. [^6] We have assumed the system is centered about 0 but that is not an issue since the long term average can be computed without knowing the ordering and subtracted from the observations.
Remark 1
If we are given an ordering of $x_{t}$s (correct or with errors) as $(x_{t_{0}},x_{t_{1}},...,x_{t_{m}})$ then we may compute covariance matrix of increments, with respect to that ordering, as the average $\frac{1}{m}\sum_{k=1}^{m}(x_{t_{k}}-x_{t_{k-1}})^{T}(x_{t_{k}}-x_{t_{k-1}})$. If $t_{0}\leq t_{1}..\leq t_{m}$ then $AA^{T}$ is the empirical covariance matrix of increments and is closely approximated by the average, by Sanov type concentration theorems (See dz). Moreover, being an average, this matrix is fairly robust to noise and small errors; a property that is critical for the algorithms in this paper.
In general, reconstructing $B$ without ordering information for the $x_{t}$ may not be possible. However, if: (1) there exists an "approximate clock" (see Definition 2) and (2) $B$ is row sparse, i.e., most of the entries in each row of $B$ are zero or close to zero; then we can approximately reconstruct $B$ with few samples in polynomial time, by solving a convex optimization problem (Algorithm 1).
Definition 2
@@
To keep the presentation simple, we will assume that if there exists a single coordinate of x(t)x(t), [^8] Here we assume a single coordinate exists as opposed to multiple coordinates or allowing a linear combination of coordinates denoted by $\tau(t)$, and it is modeled by the diffusion:
@@
where $W^{\prime}_{t}$ is standard Brownian motion and $\delta$ is positive and much greater than $\varepsilon^{2}$; then $x(t)$ is said to admit an approximate clock. We use $x_{t}$ and $x(t)$ interchangeably.
@@
We will use the approximate clock to recover an approximation of the matrix $A$, which we denote by $\tilde{A}$. We won’t actually need the ordering from approximate clock except to compute $\tilde{A}$, which gives the approximate covariance of the increments: $\tilde{A}\tilde{A}^{T}$. In Theorem 0.A.1 (stated with proof in the supplement), we formally show that ordering the $x_{i}$ values using $\tau(t)$ and computing the covariance from the increments ($\tilde{A}\tilde{A}^{T}$) closely approximates the covariance of the increments ($AA^{T}$) computed using the ordering based on the actual time.
2.1 Algorithm to reconstruct dynamical system
Algorithm 1 (below) reconstructs a dynamical system given an approximate clock. It runs in polynomial time, since the main steps are computing a covariance matrix and solving the convex optimization problem in Step 8 of Algorithm 1, which is a second order cone program, and a slight variation of basis pursuit in the compressed sensing literature (see for example rauhut). Therefore unique reconstruction is guaranteed under sparsity assumptions. In particular the following uniqueness theorem holds true for the matrix $\tilde{B}$ recovered by Algorithm 1.
Theorem 2.1
dirk Suppose each row of $B$ has at most $s$ non-zero coordinates, then for $m=\Omega(\log n)$ large and the sample times $\{t_{1},...,t_{m}\}$ for samples $\{x_{t_{1}},...,x_{t_{m}}\}$ are uniformly distributed, then with probability 1o(1)1-o(1) (for large $n$): $\|\tilde{B}-B\|_{1}=0$, i.e., $B$ can be recovered uniquely by solving a convex program. If, however, the rows are not exactly $s$ sparse and the $\ell_{1}$ norm of the nsn-s remaining entries in any row is at most $\epsilon$, then we have: $\|\tilde{B}-B\|_{1}\leq\epsilon n$.
While the proof follows by easy modifications of known results in existing compressed sensing literature (see rauhut or dirk), we note the one novelty here over existing compressed sensing literature: the formulation of the constraint in Step 8 of Algorithm 1 uses the characterization of the long term variance of the OU process (see oks).
Algorithm 1 reconstructs an approximation of the drift matrix $B$ from the covariance matrix $\Sigma$ and the (approximate) covariance matrix of the increments $\tilde{A}\tilde{A}^{T}$ up to an additive factor. It is sufficient to use $\tilde{A}\tilde{A}^{T}$ instead of $AA^{T}$ as is justified by Theorems 2.1 and 0.A.1.
@@ 2.1 Algorithm to reconstruct dynamical system
[…]
8:Solve $\tilde{B}:=\arg\min_{B}\sum_{i=0}^{n}\|B(i,\cdot)\|_{2}$   s.t. $\Sigma B+B^{T}\Sigma=-\frac{\tilde{A}\tilde{A}^{T}}{2}$
9:return $\tilde{B}$
Remark 2
@@
[…]
@@
[…]
@@
2.2 Algorithm for predicting genomic pathways
@@
1. If the gene corresponding to the terminal node of $p$ positively correlates with the pathological phenotype then it should be up-regulated by the gene preceding it in the path $p$; moreover, the same property or property (2) should hold for the subpath terminating in the penultimate gene, and so on.
2. Otherwise, if the gene corresponding to the terminal node of $p$ negatively correlates with the pathological phenotype then it should be down-regulated by the gene preceding it in the path $p$; moreover, the same property or property (1) should now hold for the subpath terminating in the penultimate gene, and so on.
@@
[…]
4:Compute and sort the list of genes in descending order of the absolute value of their correlation coefficient with the pathological phenotype, denote this list as $\vec{g}$
[…]
@@
Finally, we explain the steps required to generate pathways in Figure 1(c,d) using Algorithms 12 and 3 and RNA-seq data in the breast cancer dataset metabric; metabric2; metabric3 (the colorectal cancer dataset rc is similar).
@@
    First, we use Algorithm 1 to compute the underlying linear dynamical system matrix: $\tilde{B}$ (Figure 5(a) and Figure 6(a))
@@
    Second, we recover the perturbations $\tilde{P}$ from $\tilde{B}$ using Algorithm 2 (Figure 5(b)and Figure 6(b))
  3. 3.
    Finally, we use Algorithm 3 with $\theta=0.05$ to reconstruct the pathways similar to those in Figure 1(c,d).